# A Survey of Model Architectures in Information Retrieval

### v2

@@ 8.2 The Shifting Paradigm: From Search for Humans to Retrieval for AI
The integration of IR systems into other research domains presents new challenges. We discuss key implications for future IR modeling research as the primary "user" of retrieval shifts from humans to AI models.
@@
  • Traditional IR metrics, which are designed for human-centric evaluation, may not align with downstream task performance in retrieval-augmented systems (Lewis et al., 2020b; Petroni et al., 2021; Asai et al., 2024b). Future IR models should support end-to-end system optimization rather than focusing solely on ranking metrics (OpenAI, 2025; Huang et al., 2025).
@@
Complex tasks often require retrieving long-tail knowledge using lengthy, complex queries (Soudani et al., 2024; Su et al., 2024), demanding retrieval models capable of instruction following (Weller et al., 2024a; Ravfogel et al., 2024) and reasoning (Su et al., 2024; Shao et al., 2025; Liang et al., 2025). This has spurred the development of autonomous search agents. Existing attempts can be divided into two main directions. One line of work focuses on training more capable retrievers and rerankers. This involves creating new data pipelines to synthesize instruction-following and reasoning-focused training data (Oh et al., 2024; Weller et al., 2024b; Shao et al., 2025, inter alia) and leveraging strong backbone language models such as the proprietary GPT-4 family models (OpenAI, 2023) and open-weight Qwen (Yang et al., 2025a) to imbue the retriever with reasoning capabilities.
Another line of work treats the search/retrieval system as a static tool and focuses on improving a large reasoning model’s (LRM) ability to use it. In this setup, the LRM decides when, where, and how to conduct a search, and the results subsequently influence its further reasoning and decision-making (Nakano et al., 2021; He et al., 2025). This line of work, commonly referred to as Deep Research Agents, represents a paradigm shift from retrieval as the end goal to retrieval as tools for LLMs. Formally, we use the definition from Huang et al. (2025): deep research agents refer to AI agents powered by LLMs, integrating dynamic reasoning, adaptive planning, multi-iteration external data retrieval and tool use, and comprehensive analytical report generation for informational research tasks.
@@
Both formulations have strengths and weaknesses. API-based retrieval, in line with traditional information retrieval literature, is a fast, efficient, structured and scalable method to enable deep research agents’ access to external knowledge (Schütze et al., 2008; Singh et al., 2025). Browser-based retrieval (Nakano et al., 2021; OpenAI, 2025) has strengths of simulating real-time, interactive information seeking behaviors, in principle similar to human users’ information seeking, but incur additional latency, token consumption and introduces extra complexity encountered in the web-browsing environment. Given the current landscape of retrieval’s integration with deep research agents, an open question stands out as designing a hybrid retrieval architecture the combine the efficacy of both methods to achieve a better performance-efficiency tradeoff.
@@
From the model training perspective, recent efforts have abstracted away the details of retrieval models, treating retrieval as static tools and instead focusing on improving the capabilities of LLM-based search agents. For example, many recent efforts aim to train LRMs to use search tools more effectively via reinforcement learning or specialized fine-tuning (Li et al., 2025a; Jin et al., 2025; Li et al., 2025b; Chen et al., 2025; Wu et al., 2025b; a, inter alia). This approach is central to agentic frameworks that orchestrate tool use for complex task completion (Wu et al., 2023a; Shinn et al., 2023; Asai et al., 2024a). Despite this exciting progress, key limitations remain. To enable retrievers’ reasoning capability often requires strong backbone models (e.g., 7B scale), which can be infeasible for production systems. Even larger models (e.g., 32B scale) augmented with retrieval and trained via expensive reinforcement learning (Jin et al., 2025; Chen et al., 2025) still sometimes underperform simpler baselines with query decomposition and chain-of-thought prompting (Khot et al., 2023; Trivedi et al., 2023). A key open question is how to endow retrievers with strong reasoning capabilities using lightweight, scalable models. Another challenge lies in the joint optimization of retrievers and language models within a unified, reasoning-aware framework. Lastly, the human factors of applying such autonomous search agents remain to be studied.
@@ 8.3 Deployment, Robustness, and Trustworthiness
Traditional retrieval systems face significant challenges when scaling to web-scale document corpus, and deploying such systems requires a blend of science and engineering expertise (Dean et al., 2009; Huang et al., 2020; Li et al., 2021). In recent years, retrieval-augmented generation, conversational search, and agentic systems with memory have been widely adopted for information access (Guu et al., 2020; Lewis et al., 2020c; Google, 2019; OpenAI, 2024; Google, 2024, inter alia). These applications often require multiple rounds of retrieval and operate on dynamic corpuses, urging for efficient and effective retrieval. Mainstream inference optimization frameworks such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) have provided support for embedding models. From the modeling perspective, an open question is to design and pre-train models explicitly for retrieval purposes (Warner et al., 2024; Nussbaum et al., 2025; Günther et al., 2023).
@@
  • Robustness to AI-generated content. With the advent of LLMs, the amount of AI-generated content is also increasing. Dai et al. (2024) show that neural retrievers are biased towards AI-generated documents. Xu et al. (2024a) show that similar problems persist in text-image retrieval models. Future IR modeling research should also consider the robustness of models to AI-generated content.
@@
  • Robustness to bias and toxicity. As noted by a recent work (An et al., 2025), documents that contain biases and toxic materials can potentially jailbreak aligned LLMs. This observation highlights the importance for IR models to be robust to bias and toxic content.
@@
  • Robustness to out-of-distribution input. Given the fact that modern neural retrieval models are trained with data-driven approaches, perhaps it is not surprising to find their performance may vary with different linguistic properties of the queries and documents, i.e., out-of-distribution input from the training data. Several works have reported cross-encoder rerankers’ performance drops on out-of-domain datasets Mokrii et al. (2021); Thakur et al. (2021). In the context of retrieval-augmented generation, Cao et al. (2025) conducts a rigorous benchmarking, and finds that formality, readability, politeness and grammatical correctness — fundamental aspects of real-world user-LLM queries — can lead to significant performance variances of retrievers and RAG systems. This observation highlights the importance of retrieval models’ robustness to OOD input (Gupta et al., 2024).
We refer readers for more detailed discussions on IR models’ robustness to dedicated surveys (Asai et al., 2024b; Liu et al., 2025; Zhou et al., 2024). Addressing these robustness issues at the model architecture level is a critical and underexplored direction for future research.
@@ 9 Conclusions and Closing Thoughts
The journey of information retrieval model architectures, as we have charted, is a story of escalating abstraction and semantic depth. Beginning with the foundational principles of term-based matching in Boolean, vector space, and probabilistic models, the field systematically evolved. Learning-to-Rank introduced the power of machine learning to combine diverse statistical features, but it was the advent of neural networks that marked the first major leap toward semantic understanding. These neural ranking models, with their ability to learn representations directly from text, began to bridge the lexical gap that had long constrained traditional methods. The arrival of pre-trained transformers like BERT then catalyzed a paradigm shift, providing a powerful, universal foundation for both highly-effective cross-encoder rerankers and efficient bi-encoder retrievers. Most recently, the ascent of Large Language Models has not only scaled up these existing architectures but has also introduced entirely new paradigms, such as generative retrieval and zero-shot listwise reranking, fundamentally reshaping what a retrieval system can do.
Throughout this evolution, a core architectural tension has persisted: the trade-off between effectiveness and efficiency. This is the fundamental conflict between deep, fine-grained interaction models (like interaction-based neural rankers and cross-encoders) that offer high accuracy, and scalable representation-based models (like vector space models, dense retrievers) that enable fast, pre-computable search over massive collections. The enduring "retrieve-then-rerank" pipeline is a direct architectural answer to this trade-off. Innovations like multi-vector models (e.g., ColBERT) and hybrid sparse-dense systems represent sophisticated attempts to find a better balance on this spectrum, pushing the Pareto frontier of what is possible.
Today, we stand at another inflection point. The primary “user” of information retrieval is shifting from a human at a search bar to an AI model within a larger system. IR is no longer just a tool for finding documents; it is becoming a critical cognitive component for retrieval-augmented generation, autonomous agents, and complex reasoning systems. This shift, as outlined in §˜8, forces us to re-evaluate our core assumptions. Relevance is no longer solely about satisfying human information needs but about providing the precise factual or contextual information an AI needs to complete a downstream task. This demands new model architectures that are not only powerful and efficient but also instruction-aware, contextually flexible, and seamlessly integrable into end-to-end differentiable systems.

### v3

@@ 9.2 The Shifting Paradigm: From Search for Humans to Retrieval for AI
The integration of IR systems into other research domains presents new challenges. We discuss key implications for future IR modeling research as the primary “user” of retrieval shifts from humans to AI models.
@@
  • Traditional IR metrics, which are designed for human-centric evaluation, may not align with downstream task performance in retrieval-augmented systems (Lewis et al., 2020b; Petroni et al., 2021; Asai et al., 2024c). Future IR models should support end-to-end system optimization rather than focusing solely on ranking metrics (OpenAI, 2025; Huang et al., 2025).
Retrieval Augmented Generation.
Retrieval Augmented Generation (RAG) refers to architectures that combine an external retrieval module with a generative model, enabling the system to access and condition on external knowledge before producing output. Early RAG formulations separate retrieval and generation as distinct components: a dense or sparse retriever extracts relevant passages for a given input, and a generator (usually an LLM) conditions on those retrieved results to produce text (Lewis et al., 2020b; Ram et al., 2023b; Mallen et al., 2023; Asai et al., 2024c; Chen et al., 2020b; Xu & Cohen, 2023). This architectural decoupling allows retrieval to be optimized independently from generation, which is useful for modularity and scalability. A typical RAG pipeline comprises:
  • Corpus embedding and indexing (vector stores or sparse indexes) for fast retrieval;
  • Retriever model (dense vectors or hybrid techniques) that returns top-$k$ relevant items;
  • Fusion mechanism that combines retrieved content with the query (e.g., concatenation or cross-attention) before passing it to the generator;
  • Generator that produces the final answer based on fused context.
This simple architecture improves factual accuracy and reduces hallucination by grounding generation on external context. However, it also introduces key engineering and modeling considerations:
  • Retriever-Generator Coupling: The architectural choice of where and how to combine retrieval results can affect quality and efficiency. Early approaches rely on simple concatenation of text passages into the generator’s context window (Lewis et al., 2020b; Ram et al., 2023b), which do not alter the architecture of the generator. More advanced designs use cross-attention layers or rerankers to improve precision (Izacard & Grave, 2021; Park et al., 2023; An et al., 2025b).
  • Multi-Modal and Structured Retrieval: Recent work explores RAG variants that extend beyond plain text. For example, GraphRAG and related graph-enhanced frameworks integrate structured knowledge sources, enabling retrieval of relational paths not easily captured by vector similarity alone (Edge et al., 2025; Li et al., 2025a). These architectures introduce additional components like graph encoders and reasoning layers, allowing the system to extract multi-hop or entity-centric context before generation.
  • Adaptive Retrieval: Newer RAG pipelines incorporate reasoning capabilities into the generator, to dynamically decide whether to retrieval or to generate answer (Trivedi et al., 2023; Jiang et al., 2023b; Asai et al., 2024b).
These architectural choices have direct implications: strong retrieval can reduce generation errors but may increase pipeline complexity and latency; graph-aware models can support structured reasoning at the cost of larger indexes and additional components; adaptive retrieval strategies can improve effectiveness but require careful task planning and retrieval coordination.
Overall, RAG architectures represent a spectrum from simple retrieval-plus-generation pipelines to complex multi-component systems that integrate heterogeneous knowledge representations and adaptive retrieval planning to improve both relevance and reasoning capability. We refer readers to (Gao et al., 2023; Li et al., 2025e; Mei et al., 2025) for more comprehensive reviews of this topic.
@@
Complex tasks often require retrieving long-tail knowledge using lengthy, complex queries (Soudani et al., 2024; Su et al., 2024), demanding retrieval models capable of instruction following (Weller et al., 2025a; Ravfogel et al., 2024) and reasoning (Su et al., 2024; Shao et al., 2025; Liang et al., 2025). This has spurred the development of autonomous search agents. Existing attempts can be divided into two main directions. One line of work focuses on training more capable retrievers and rerankers. This involves creating new data pipelines to synthesize instruction-following and reasoning-focused training data (Oh et al., 2024; Weller et al., 2024; Shao et al., 2025, inter alia) and leveraging strong backbone language models such as the proprietary GPT-4 family models (OpenAI et al., 2023) and open-weight Qwen (Yang et al., 2025a) to imbue the retriever with reasoning capabilities.
Another line of work treats the search/retrieval system as a static tool and focuses on improving a large reasoning model’s (LRM) ability to use the retrieval tool to improve reasoning and handle in-context memory (Jin et al., 2025; Yu et al., 2025; Gao et al., 2025; Xu et al., 2025d, inter alia). In this setup, the LRM decides when, where, and how to conduct a search, and the results subsequently influence its further reasoning and decision-making (Nakano et al., 2021; He et al., 2025). This line of work, commonly referred to as Deep Research Agents, represents a paradigm shift from retrieval as the end goal to retrieval as tools for LLMs. Formally, we use the definition from Huang et al. (2025): deep research agents refer to AI agents powered by LLMs, integrating dynamic reasoning, adaptive planning, multi-iteration external data retrieval and tool use, and comprehensive analytical report generation for informational research tasks.
@@
Both formulations have strengths and weaknesses. API-based retrieval, in line with traditional IR literature, is a fast, efficient, structured and scalable method to enable deep research agents’ access to external knowledge (Schütze et al., 2008; Singh et al., 2025b). Browser-based retrieval (Nakano et al., 2021; OpenAI, 2025) offers the advantage of simulating real-time, interactive information-seeking behavior — conceptually similar to that of human users. However, it incurs additional latency and token consumption, and introduces further complexity arising from the web-browsing environment. Given the current landscape of retrieval integration within deep research agents, a central open question is how to design a hybrid retrieval architecture that combines the strengths of both paradigms to achieve a superior performance-efficiency tradeoff.
@@
From the model training perspective, recent efforts have abstracted away the details of retrieval models, treating retrieval as static tools and instead focusing on improving the capabilities of LLM-based search agents. For example, many recent efforts aim to train LRMs to use search tools more effectively via reinforcement learning or specialized fine-tuning (Li et al., 2025b; Jin et al., 2025; Li et al., 2025c; Chen et al., 2025; Wu et al., 2025b; a, inter alia). This approach is central to agentic frameworks that orchestrate tool use for complex task completion (Wu et al., 2023a; Shinn et al., 2023; Asai et al., 2024b). Despite this exciting progress, key limitations remain. To enable retrievers’ reasoning capability often requires strong backbone models (e.g., 7B scale), which can be infeasible for production systems. Even larger models (e.g., 32B scale) augmented with retrieval and trained via expensive reinforcement learning (Jin et al., 2025; Chen et al., 2025) still sometimes underperform simpler baselines with query decomposition and chain-of-thought prompting (Khot et al., 2023; Trivedi et al., 2023). A key open question is how to endow retrievers with strong reasoning capabilities using lightweight, scalable models. Another challenge lies in the joint optimization of retrievers and language models within a unified, reasoning-aware framework. Lastly, the human factors of applying such autonomous search agents remain to be studied. We refer readers to (Singh et al., 2025a; Liang et al., 2025; Xi et al., 2025; Lin et al., 2025) for more comprehensive reviews of this topic.
9.3 Retrieval Beyond Simple Relevance: Instruction-Following and Reasoning-Aware Retrieval
Instruction-Following Retrieval.
Instruction-following retrieval extends the standard query-document formulation by conditioning the retriever on a natural-language instruction. Architecturally, this is typically implemented by jointly encoding the instruction and the query in a shared bi-encoder, for example through simple concatenation or lightweight attention layers, so that the query representation adapts to task intent (Weller et al., 2025a; Oh et al., 2024). Training objectives often use contrastive learning with instruction-conditioned positives and hard negatives, allowing the model to capture fine-grained task semantics without large cross-encoder computations (Weller et al., 2024). Instruction-aware retrievers benefit from curated or synthetic datasets pairing instructions, queries, and relevant passages, which improves robustness to paraphrasing and query phrasing variations. These design choices preserve scalability while improving performance on instruction-heavy tasks. Hybrid pipelines may further combine instruction-conditioned retrieval with downstream rerankers or fusion mechanisms, ensuring that retrieved evidence aligns with the specific needs of generation or reasoning tasks (Shao et al., 2024; Weller et al., 2025b).
Reasoning-Aware Retrieval.
Reasoning-aware retrieval aims to retrieve evidence that supports multi-step inference and complex reasoning (Su et al., 2024). One line of work uses graph-based architectures, where the retriever constructs and encodes relational graphs or multi-hop chains of passages. These components — graph construction, encoding, path selection, and ranking — allow the retriever to return connected evidence that reflects reasoning dependencies (Liu et al., 2025a; Edge et al., 2025). While effective for compositional queries, these approaches increase system complexity, latency, and index size.
A different line of work, exemplified by ReasonIR (Shao et al., 2025) and RankR1 (Zhuang et al., 2025), improves reasoning capability through data curation and task-aligned training rather than altering the underlying “retrieve-then-rerank” pipeline. By generating multi-step reasoning datasets or instruction-guided examples, these methods teach standard retrievers and rerankers to prioritize evidence that is useful for downstream reasoning, achieving stronger performance without introducing graph-based components. From an evaluation perspective, reasoning-aware retrieval highlights the need for task-centered metrics that measure downstream reasoning success, emphasizing the tight connection between architectural or training choices and reasoning effectiveness.
@@ 9.4 Deployment, Robustness, and Trustworthiness
Traditional retrieval systems face significant challenges when scaling to web-scale document corpus, and deploying such systems requires a blend of science and engineering expertise (Dean, 2009; Huang et al., 2020; Li et al., 2021). In recent years, retrieval-augmented generation, conversational search, and agentic systems with memory have been widely adopted for information access (Guu et al., 2020; Lewis et al., 2020b; Nayak, 2019; OpenAI, 2024; Google, 2024, inter alia). These applications often require multiple rounds of retrieval and operate on dynamic corpuses, urging for efficient and effective retrieval. Mainstream inference optimization frameworks such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024b) have provided support for embedding models. From the modeling perspective, an open question is to design and pre-train models explicitly for retrieval purposes (Warner et al., 2025; Nussbaum et al., 2025; Günther et al., 2024).
@@
  • Robustness to AI-generated content. With the advent of LLMs, the amount of AI-generated content is also increasing. Dai et al. (2024) showed that neural retrievers are biased towards AI-generated documents. Xu et al. (2024a) showed that similar problems persist in text-image retrieval models. Future IR modeling research should also consider the robustness of models to AI-generated content.
@@
  • Robustness to bias and toxicity. As noted by a recent work (An et al., 2025a), documents that contain biases and toxic materials can potentially jailbreak aligned LLMs. This observation highlights the importance for IR models to be robust to bias and toxic content.
@@
  • Robustness to out-of-distribution input. Given the fact that modern neural retrieval models are trained with data-driven approaches, perhaps it is not surprising to find their performance may vary with different linguistic properties of the queries and documents, i.e., out-of-distribution input from the training data. Several works have reported cross-encoder rerankers’ performance drops on out-of-domain datasets (Mokrii et al., 2021; Thakur et al., 2021). In the context of retrieval-augmented generation, Cao et al. (2025) conducted a rigorous benchmarking, and found that formality, readability, politeness and grammatical correctness — fundamental aspects of real-world user-LLM queries — can lead to significant performance variances of retrievers and RAG systems. This observation highlights the importance of retrieval models’ robustness to out-of-distribution (OOD) input (Gupta et al., 2024a).
We refer readers for more detailed discussions on IR models’ robustness to dedicated surveys (Asai et al., 2024c; Liu et al., 2025c; Zhou et al., 2024). Addressing these robustness issues at the model architecture level is a critical and underexplored direction for future research.
@@ 10 Conclusions and Closing Thoughts
The journey of IR model architectures, as we have charted, is a story of escalating abstraction and semantic depth. Beginning with the foundational principles of term-based matching in Boolean, vector space, and probabilistic models, the field systematically evolved. Learning-to-Rank introduced the power of machine learning to combine diverse statistical features, but it was the advent of neural networks that marked the first major leap toward semantic understanding. These neural ranking models, with their ability to learn representations directly from text, began to bridge the lexical gap that had long constrained traditional methods. The arrival of pre-trained transformers like BERT then catalyzed a paradigm shift, providing a powerful, universal foundation for both highly-effective cross-encoder rerankers and efficient bi-encoder retrievers. Most recently, the ascent of LLMs has not only scaled up these existing architectures but has also introduced entirely new paradigms, such as generative retrieval and zero-shot listwise reranking, fundamentally reshaping what a retrieval system can do.
Throughout this evolution, a core architectural tension has persisted: the tradeoff between effectiveness and efficiency. This is the fundamental conflict between deep, fine-grained interaction models (like interaction-based neural rankers and cross-encoders) that offer high accuracy, and scalable representation-based models (like vector space models, dense retrievers) that enable fast, pre-computable search over massive collections. The enduring “retrieve-then-rerank” pipeline is a direct architectural answer to this tradeoff. Innovations like multi-vector models (e.g., ColBERT) and hybrid sparse-dense systems represent sophisticated attempts to find a better balance on this spectrum, pushing the Pareto frontier of what is possible.
Today, we stand at another inflection point. The primary “user” of IR is shifting from a human at a search bar to an AI model within a larger system. IR is no longer just a tool for finding documents; it is becoming a critical cognitive component for retrieval-augmented generation, autonomous agents, and complex reasoning systems. This shift, as outlined in Section˜9, forces us to re-evaluate our core assumptions. Relevance is no longer solely about satisfying human information needs but about providing the precise factual or contextual information an AI needs to complete a downstream task. This demands new model architectures that are not only powerful and efficient but also instruction-aware, contextually flexible, and seamlessly integrable into end-to-end differentiable systems.
@@
Acknowledgments
This work was partially supported by grants from NSF (DMS 2134223 and IIS 2205418).